# Low loss monolithic barium titanate on insulator integrated photonics with intrinsic quality factor million

### v1

@@ 2 Department of Mechanical Science and Engineering,
0.3in-0.5in Publication Year Device Type Platform Intrinsic Q-factor () Loaded Q-factor () Propagation loss (, dB/cm) Wavelength (nm) This work Racetrack Resonator BTOI 1.35 x 106 0.80 x 106 0.32 1520 Ring Resonator BTOI 0.85 x 106 0.59 x 106 0.52 1520 Isti et. al. [isti_fabrication_2025] 2025 Waveguide BTOI N/A N/A 0.7 1546 Raju et. al. [raju_highq_2025] 2025 Racetrack Resonator BTOI 0.5 x 106 0.09 x 106 0.62† 1550 Ring Resonator BTOI 0.2 x 106 0.10 x 106 2.11 1550 Riedhauser et. al. [riedhauser_absorption_2025] 2025 Resonator BTOI 0.45 x 106 N/A 0.86 1554 Möhl et. al. [mohl_bidirectional_2025] 2025 Resonator BTOI 0.25 x 106† ( = 776 MHz at 1550 nm) N/A N/A 1537–1577 Dong et. al. [dong_monolithic_2023] 2023 Waveguide BTOI N/A N/A 3.17 1535–1565 Petraru et. al. [petraru_ferroelectric_2002] 2002 Waveguide BTOI N/A N/A 2–3 632 Tao et. al. [tao_towards_2024] 2023 Waveguide Si/BTO N/A N/A 1 1545 Posadas et. al. [posadas_rf-sputtered_2023] 2023 Waveguide Si/BTO N/A N/A 1.08 1550 Abel et. al. [abel_large_2019] 2019 Waveguide Si/BTO N/A N/A 10 1541 Eltes et. al. [eltes_batio3-based_2019] 2019 Waveguide Si/BTO N/A N/A 3 1550–1551 Resonator Si/BTO N/A 0.05 x 106 5.8 Abel et. al. [abel_hybrid_2016] 2016 Resonator Si/BTO N/A 0.0046 x 106 47 1555–1575 Eltes et. al. [eltes_low-loss_2016] 2016 Waveguide Si/BTO N/A N/A 6 1550 Kohli et. al. [kohli_plasmonic_2023] 2023 Waveguide (plasmonic) SiN/BTO N/A N/A 5000 1557 Eltes et. al. [eltes_integrated_2020] 2020 Waveguide SiN/BTO N/A N/A 5.6 1511–1512 Ortmann et. al. [ortmann_ultra-low-power_2019] 2019 Resonator SiN/BTO N/A N/A 9.4 1550

### v2

@@ 2 Department of Mechanical Science and Engineering,
| Publication | Year | Device Type | Platform | Intrinsic Q-factor () | Loaded Q-factor () | Propagation loss (, dB/cm) | Wavelength (nm) |
| This work | Racetrack Resonator | BTOI | 1.35 x 106 | 0.80 x 106 | 0.32 | 1520 | |
| Ring Resonator | BTOI | 0.85 x 106 | 0.59 x 106 | 0.52 | 1520 | ||
| Isti et. al. [isti_fabrication_2025] | 2025 | Waveguide | BTOI | N/A | N/A | 0.7 | 1546 |
| Raju et. al. [raju_highq_2025] | 2025 | Racetrack Resonator | BTOI | 0.5 x 106 | 0.09 x 106 | 0.62† | 1550 |
| Ring Resonator | BTOI | 0.2 x 106 | 0.10 x 106 | 2.11 | 1550 | ||
| Riedhauser et. al. [riedhauser_absorption_2025] | 2025 | Resonator | BTOI | 0.45 x 106 | N/A | 0.86 | 1554 |
| Möhl et. al. [mohl_bidirectional_2025] | 2025 | Resonator | BTOI | 0.25 x 106† ( = 776 MHz at 1550 nm) | N/A | N/A | 1537–1577 |
| Dong et. al. [dong_monolithic_2023] | 2023 | Waveguide | BTOI | N/A | N/A | 3.17 | 1535–1565 |
| Petraru et. al. [petraru_ferroelectric_2002] | 2002 | Waveguide | BTOI | N/A | N/A | 2–3 | 632 |
| Tao et. al. [tao_towards_2024] | 2023 | Waveguide | Si/BTO | N/A | N/A | 1 | 1545 |
| Posadas et. al. [posadas_rf-sputtered_2023] | 2023 | Waveguide | Si/BTO | N/A | N/A | 1.08 | 1550 |
| Abel et. al. [abel_large_2019] | 2019 | Waveguide | Si/BTO | N/A | N/A | 10 | 1541 |
| Eltes et. al. [eltes_batio3-based_2019] | 2019 | Waveguide | Si/BTO | N/A | N/A | 3 | 1550–1551 |
| Resonator | Si/BTO | N/A | 0.05 x 106 | 5.8 | |||
| Abel et. al. [abel_hybrid_2016] | 2016 | Resonator | Si/BTO | N/A | 0.0046 x 106 | 47 | 1555–1575 |
| Eltes et. al. [eltes_low-loss_2016] | 2016 | Waveguide | Si/BTO | N/A | N/A | 6 | 1550 |
| Kohli et. al. [kohli_plasmonic_2023] | 2023 | Waveguide (plasmonic) | SiN/BTO | N/A | N/A | 5000 | 1557 |
| Eltes et. al. [eltes_integrated_2020] | 2020 | Waveguide | SiN/BTO | N/A | N/A | 5.6 | 1511–1512 |
| Ortmann et. al. [ortmann_ultra-low-power_2019] | 2019 | Resonator | SiN/BTO | N/A | N/A | 9.4 | 1550 |